OpenStudy (anonymous):

which of the following represents x=1/2y written in general form

OpenStudy (anonymous):

are there any answers for this problem on your homework?

OpenStudy (anonymous):

like a b or c d

OpenStudy (anonymous):

a= 1/2x-y=0 b=2x-y=0 c=x-1/2y=0